OpenStudy (anonymous):

Which expression is a cube root of -2?

OpenStudy (anonymous):

\[\sqrt[3]{2}(\cos(60)+isin(60))\]

OpenStudy (anonymous):

\[\sqrt[3]{2}(\cos(120)+isin(120))\]

OpenStudy (anonymous):

\[\sqrt[3]{2}(\cos(260)+isin(260)\]

OpenStudy (anonymous):

\[\sqrt[3]{2}(\cos(90)+isin(90))\]

OpenStudy (misty1212):



OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

the real cube root of \(-2\) sits right on the real axis the other two divide the circle up in to 3 parts

OpenStudy (misty1212):

with the real cube root of two as one of them of course 3 parts because it is the cubed root it should be clear from the picture what the angles are

OpenStudy (anonymous):

okay, so the angles are 60, 180, and 300?

OpenStudy (misty1212):

yeah

OpenStudy (misty1212):

although adults use radians, not degrees, but whatever

OpenStudy (anonymous):

does that mean it's the first option?

OpenStudy (misty1212):

yes